OpenStudy (anonymous):

A druggist has two mixtures of a certain chemical and water, one containing 10% of the chemical and the other containing 40% of the chemical. How much of each mixture should he use to make 10 ounces that contain 25% of the chemical? If x is the amount of 10% mixture used, then which expression represents the ounces of chemical contained in it?

OpenStudy (anonymous):

0.1x x 10x

OpenStudy (anonymous):

10% of x is equal to 0.1x Let say I used A ounces from the first mixture and B ounces from the second one. Total of chemicals = 10*25% = A*10% + B*40% 10*0.25 = 2.5 = 0.1A+0.4B Total of mixtures = 10 = A + B You have two unknowns and two equations. Can you solve them?

OpenStudy (anonymous):

umm no lol sorry

OpenStudy (anonymous):

nvm i got it thanks

OpenStudy (anonymous):

When you have n unknows you need n equations two find them. Luckily we have 2 unknowns and 2 equations, we can solve them. First find one of the A or B (doesn't matter which one you find)

OpenStudy (anonymous):

If you want I can solve them. It's very simple.

OpenStudy (anonymous):

to late lol thanks tho can you help with another problem?

OpenStudy (anonymous):

Solve for y: y+10/y+3 = y+1/y-1